pleas help me solve the equation:
Anne flew her plane 400 km with the wind in the same time it took her to go 200 km against the wind. The speed of the wind is 50km per hour. Fine the speed of the plane in still air.

Answer by mananth(16949)

plane speed x mph
wind speed 50 mph
..
against wind x - 50 mph
with wind x + 50 mph

Distance against wind 200
Distance with wind 400
Time = same
200 /(x- 50 )= 400 /(x+ 50 )
200 *(x+ 50 )= 400 *(x- 50 )
200 x + 10000 = 400 x -20000
400 x -200 x = 20000 + 10000
200 x = 30000
/ 200
x= 150 mph